a) Suppose you have borrowed Rs 10,00,000 from your employer for 3 years at 9% rate of interest to buy a car. If your employer requires you to pay back in 3 equal end of year instalments, prepare the loan amortisation schedule.
b) Assume to finance your post graduate studies in an executive weekend programme, you undertake a part time job for 5 years. Your employer fixes your annual salary at Rs 10,00,000 with a provision that you will get an annual increment of 10%. If you will save it all and be able to make your investment grow @ 12% p.a., calculate the present value of your salary? What amount will be able to accumulate by the end of 5th year?
Answer #1
Step-1:Calculation of annual payment
Annual payment = Loan amount / Present value of annuity of 1
= ₹ 10,00,000.00 / 2.531294666
= ₹    3,95,054.76
Working:
Present value of annuity of 1 = (1-(1+i)^-n)/i Where,
= (1-(1+0.09)^-3)/0.09 i = 9%
= 2.531294666 n = 3
Step-1:Calculation of amortization Schedule
Year Beginning loan Interest Expense Annual Payment Reduction of Liability Ending Loan
a b=a*9% c d=c-b e=a-d
1 ₹ 10,00,000.00 ₹        90,000.00 ₹    3,95,054.76 ₹    3,05,054.76 ₹ 6,94,945.24
2 ₹    6,94,945.24 ₹        62,545.07 ₹    3,95,054.76 ₹    3,32,509.69 ₹ 3,62,435.56
3 ₹    3,62,435.56 ₹        32,619.20 ₹    3,95,054.76 ₹    3,62,435.56 ₹                0.00
Total ₹    1,85,164.27 ₹ 11,85,164.27 ₹ 10,00,000.00